OpenStudy (anonymous):

Find the equation(s) of the tangent line(s) to the parabola f(x)=x^2 through the given points (not on the graph). A) (0,a) B) (a,0)

OpenStudy (anonymous):

you know that the slope of any point on the graph of \((x,x^2)\) is \(2x\) and the slope through \((x,x^2)\) and \((0,a)\) is \(\frac{x^2-a}{x-0}=\frac{x^2-a}{x}\) set \[\frac{x^2-a}{x}=2x\] and solve for \(x\) second one is similar

OpenStudy (anonymous):

Thanks!